Find the sum of the digits of the number (10^2007 + 5)/3
A. 6018. B. 6020. C. 6023. D. 6026. E. None of these

Found 2 solutions by CharlesG2, Alan3354:
Answer by CharlesG2(834)

Find the sum of the digits of the number (10^2007 + 5)/3
A. 6018. B. 6020. C. 6023. D. 6026. E. None of these
this number will have 2006 digits, 2005 of them will be 3's, last one will be a 5
2005*3 = 6015
6015+5 = 6020
answer is B



Answer by Alan3354(69443)

Find the sum of the digits of the number (10^2007 + 5)/3
A. 6018. B. 6020. C. 6023. D. 6026. E. None of these
--------------------------
10^2007 = 1 and 2007 zeroes.
Adding 5 and dividing by 3 gives 2006 3's and a 5
--> 2006*3 + 5 = 6023